3. Gauss Lau: Weather balloons are typically filled with hydrogen and reach a diameter of 8 metres in the upper atmosphere.
Answer #1

discharged I hydrogen nons. r (rum) Gaussian surface as consides any spherical surface and thickness dae do = 40x²deh of radi=> 4 4 [de em Pu 4992 do = 10-12x40 4 25 S 0 = 0.5k X10 12 x 1024. E 34008² = Sin E=9x10xo. Sax1024x10-12 82 E = 23150:592X10